OpenStudy (anonymous):

find vertex of parabola f(x)= 5-(x+20) squared. [ the (x+20) is squared. I don't know how to write that]

OpenStudy (benjieg6pd):

The vertex of a parabola in the form y = a(x - h)^2 + k is the point with coordinates (h , k) The parabola f(x)=5 - (x + 20)^2 can be written in the above format as: y = -(x + 20)^2 + 5 Or y = -1(x - -20)^2 + 5. In this format you can see the values of the coefficients a=-1, h=-20, k=5. So the vertex is the point with coordinates ( -20, 5)